OpenStudy (anonymous):

Solve by factoring: 3x2 + 2x − 1 = 0

OpenStudy (anonymous):

(3x-1)(x+1)

OpenStudy (anonymous):

3x2 + 2x − 1 = 0 3x2 + 3x -x− 1 = 0 (3x-1)(x+1)